Rea is 20 years older than Tina. Sixteen years ago, Rea was three times as old as Tina was then. How old are Rea and Tina now?

Found 2 solutions by ankor@dixie-net.com, ikleyn:
Answer by ankor@dixie-net.com(22740)

let r = Rea's age now
let t = Tina's
:
Write an equation for each statement, simplify as much as possible.
:
Rea is 20 years older than Tina.
r = t + 20
Sixteen years ago, Rea was three times as old as Tina was then.
r - 16 = 3(t-16)
r - 16 = 3t - 48
r = 3t - 48 + 16
r = 3t - 32
How old are Rea and Tina now?
replace r with (t+20)
t + 20 = 3t - 32
20 + 32 = 3t - t
52 = 2t
t = 52/2
t = 26 yrs is Tina's age
then
r = 26 + 20
r = 46 yrs is Rea's age
:
;
:
Check this in the statement:
"Sixteen years ago, Rea was three times as old as Tina was then."
46 - 16 = 3(26 - 16)
30 = 3(10)
:

Answer by ikleyn(52787)

.
There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problem for the day of April, 1
    - OVERVIEW of lessons on age problems
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

It contains a lot of other word problems and many other interesting and useful things.